Capital Flows & Demography

Dave Backus

Based on work by Tom Cooley, Espen Henriksen, and Jan Möller

Canadian Macroeconomic Study Group Montreal | November 2, 2012

This version: November 8, 2012

Backus (NYU)  Capital Flows & Demography  1 / 35

Outline

Capital flows: thoughts Capital flows: evidence Demography Model (work in progress) Saving and investment in China

Thoughts

Capital Flows: Thoughts

Backus (NYU)  Capital Flows & Demography  2 / 35

Thoughts

Thoughts about capital flows

Michael Bordo, “Globalization in historical perspective,” 2002

The fifty years before World War I saw massive flows of capital from Western Europe to (mainly) the Americas and Australasia. At its peak, the outflow from Britain reached nine percent of GNP and was almost as high in France, Germany, and the Netherlands. Private capital moved essentially without restriction ... into bonds that financed railroads and other infrastructure — and into government debt.

Link: http://econweb.rutgers.edu/bordo/nabe.pdf

Thoughts about capital flows

Lars Jonung, “Sweden under the gold standard,” 1984

In the second half of the nineteenth century, Sweden began a process of industrialization financed by considerable capital imports which transformed a basically agrarian country into an industrialized society.

From Bordo & Schwartz, Retrospective on the Gold Standard

Thoughts

Thoughts about capital flows Thoughts about capital flows

Michael Bordo, “Globalization in historical perspective,” 2002

A striking feature is the size and persistence of current account deficits in this period, esp in Australia, Canada, Argentina, and the Nordic countries, as well as the surpluses of the UK and France.

This globalization of finance had its dark side — in periodic crises when capital flows abruptly reversed themselves.

Link: http://econweb.rutgers.edu/bordo/nabe.pdf

Thoughts about capital flows

John Maynard Keynes, 1941-42

I disagree most strongly that the control of capital movements may be unnecessary. ... I see no reason to feel confidence that the more stable conditions [of the post-war era] will remove the more dangerous movements [of capital].

Central control of capital movements, both inward and outward, should be a permanent feature of the post-war system.

Quoted by Crotty, JEL, 1983

Thoughts about capital flows

Daniel Gross, New York Times, May 8, 2005

[US] imbalances are eerily reminiscent of recent economic crises. Could we see a perfect storm [for the US economy]? If so, what would it look like?

⋆ Nouriel Roubini estimates that long-term interest rates in the US could rise sharply and the dollar fall.

⋆ Jeffrey Frankel adds: “some of us have been warning of this hard-landing scenario for more than 20 years.”

Thoughts about capital flows

Group of 20, Communique, April 15-16, 2011

We agreed on a set of indicative guidelines ... to address persistently large imbalances. We now launch ... an in-depth assessment of the nature of these imbalances and the root causes of impediments to adjustment. ... We will ascertain for our next meeting the corrective and preventive measures.

Øystein Olsen, Norges Bank, March 2011

Global trade imbalances have been reduced somewhat over the past two years, but there is a considerable risk that they will persist. They must be corrected.

Thoughts about capital flows

Dick Caves and Ron Jones, World Trade and Payments, 1973

By accounting rule, credits (or receipts) make the balance positive or “favorable.” Debits are “unfavorable.” Notice the gravitational pull of semantics! The receipts side owns all the good words, and has done so since the eighteenth-century mercantilists made a virtue of “storing up treasure.”

(first edition, page 87, lightly edited)

Thoughts

Thoughts: summary

Long history of ambivalence over capital flows

▶ Facilitator of economic growth? ▶ Or source of instability?

Suggestion: don’t let the words decide

▶ Find and replace:

“global imbalances” ⇒ “international capital flows”

Our focus

▶ Where do capital flows come from? ▶ Why are they so persistent?

Evidence

Capital Flows: Evidence

Backus (NYU)  Capital Flows & Demography  12 / 35

Evidence

Evidence: capital flows

Largest inflows Amount (b$US) Largest outflows Amount (b$US)
United States  –470 China  306
Italy  –72 Japan  195
Spain  –63 Germany  176
United Kingdom  –56 Mideast oil  150
France  –53 Switzerland  75
India  –49 Russia  71
Turkey  –49 Netherlands  56
Canada  –49 Norway  53
Brazil  –47 Taiwan  41
Greece  –32 Sweden  30
Australia  –32 Korea  28
Portugal  –23 Malaysia  28
Total inflows  –1183 Total outflows  1466

Source: IMF, WEO, April 2011 version, data for 2010

Backus (NYU)  Capital Flows & Demography  13 / 35

Evidence

Evidence: capital flows

Evidence

Evidence: net foreign assets

Evidence

Evidence: capital flows and stocks

Capital flow variation is low-frequency (so much for HP filtering)

Ditto net foreign assets

No evident tendency for either to move toward zero (where’s the “external balance condition”?)

Demography

Demography

Backus (NYU)  Capital Flows & Demography  17 / 35

Demography

Evidence: demography

Model

Model (a progress report)

Backus (NYU)  Capital Flows & Demography  19 / 35

Model

Model: motivation

Capital flows and stocks are persistent Demography inherently persistent — and different across countries Worth exploring a connection? [Apparently yes: Attanasio-Kitao-Violante, Bloom-Canning, Boersch-Supan-Winter, Brooks, Domeij-Floden, Feroli, Ferrero, Henriksen, Krueger-Ludwig, and others all had the same idea]

Model

Model: structure

One-good world Overlapping generations, realistic mortality rates Key ages: start working/consuming at 21, retire at 65 Preferences: power utility over consumption, fixed labor supply Technology: CES production, country-specific productivity Today: explore impact of changes in life expectancy

Model

Model: mortality

Model

Model: age distribution

Model

Model: net foreign assets

Capital allocated to equate marginal products

▶ Connected to age distribution via effective labor input ▶ Equates K/Y across countries

Net worth follows from life-cycle saving

▶ Saving and net worth depend on life expectancy ▶ Also age distribution

NFA is difference between net worth and capital stock

Model

Model: steady state calculations

Impact of life expectancy Two versions

▶ Closed economy: capital stock and interest rate adjust ▶ Open economy: interest rate fixed, NFA adjusts

Model

Model: closed economy net asset profile

Model

Model: closed economy capital and wealth

Model

Model: closed economy interest rate

Model

Model: open economy net asset profile

Model

Model: open economy capital and wealth

China

China

Backus (NYU)  Capital Flows & Demography  31 / 35

China

Why are China’s saving and investment rates so high?

China

Does China have too much capital?

China

Does China have too much capital?

China

Why is saving so high?

Demography? Precautionary saving? High “business saving”? Other?

Final thoughts

Last thoughts

Capital flows

▶ A fact of life for more than a century ▶ Persistent ▶ A role for demography? ▶ What else would you suggest?

Open questions

▶ Why are capital-output ratios so different across countries? ▶ Why are saving rates so different across countries? ▶ Why is China’s so high?